1. Profit maximization Suppose Hannah gives haircuts on Saturdays to make extra money. She is the only person in town cuttingFill in the missing cells of the table and then use them to answer the questions that follow. Total Marginal Marginal Marginagraph from left to right, starting with zero haircuts and ending with five. Line segments will automatically connect the poinMC AND MR (Dollars per week) Marginal Revenue Marginal Cost OUTPUT (Haircuts per week) Help Clear All Hannah maximizes her prHannah maximizes her profit by serving per week and charging per haircut. If Hannah gave more haircuts than her optimal quant
Answer #1

(1) When Quantity (Q) = 4,

(a) Total revenue (TR) = Price x Q = 21 x 4 = 84

(b) Marginal revenue (MR) = change in TR / change in Q = (84 - 72) / (4 - 3) = 12/1 = 12

(c) Marginal cost (MC) = change in TC / change in Q = (35 - 18) / (4 - 3) = 17/1 = 17

(d) Profit = TR - TC = 84 - 35 = 49

(e) Marginal profit = change in profit / change in Q = (49 - 54) / (4 - 3) = - 5/1 = - 5

(2)

COST, REVENUE, AND PROFIT (Dollars per week) 100 Total Revenue Total Cost Profit 2 3 4 5 OUTPUT (Haircuts per weekl Help Clea

(3)

MC AND MR (Dollars per week) Marginal Revenue O Marginal Cost 1 2 3 4 5 OUTPUT (Haircuts per week) Help Clear All

(4) Hanna maximizes profit by serving 3 haircuts (when MR > MC) by charging $24.

(5) If she gives more haircuts (Q > 3), following is true:

- Marginal profit is negative

- Marginal revenue is less than marginal cost

- Total profit will decline from optimal